A bomber on a military mission is flying at a height of 300m above the ground at 60km/mi. it drops a bomber on a Target on the ground. Determine the acute angle between the vertical angle and the line joining the bomber and the target at instant that bomb is realeased.
1
Expert's answer
2018-10-30T19:07:09-0400

t=sqrt(2h/g) = 7.7 s

l = tv = 7700 m

phi = atan(7700/300) = 87.8 deg